Find: (i) 0.2 × 6 (ii) 8 × 4.6 (iii) 2.71 × 5 (iv) 20.1 × 4 (v) 0.05 × 7 nveahuyghuyhdsbvhvuefwqtwyt8734yu85h8gueqwgy289yu38huj8ehyu

Find:
(i) 0.2 × 6

(ii) 8 × 4.6

(iii) 2.71 × 5

(iv) 20.1 × 4

(v) 0.05 × 7

 

nveahuyghuyhdsbvhvuefwqtwyt8734yu85h8gueqwgy289yu38huj8ehyu

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

(i) 0.2 × 6

Solution:- We have,

= (2/10) × 6

= (12/10)
On dividing a decimal by 10, the decimal point is shifted to the left by one place. Then,

= 1.2

(ii) 8 × 4.6

Solution:- We have,

= (8) × (46/10)

= (368/10)
On dividing a decimal by 10, the decimal point is shifted to the left by one place. Then,

= 36.8

(iii) 2.71 × 5

Solution:- We have,

= (271/100) × 5

= (1355/100)
On dividing a decimal by 100, the decimal point is shifted to the left by two places. Then,

= 13.55

(iv) 20.1 × 4

Solution:- We have,

= (201/10) × 4

= (804/10)
On dividing a decimal by 10, the decimal point is shifted to the left by one place. Then,

(v) 0.05 × 7

Solution:- We have,

= (5/100) × 7

= (35/100)
On dividing a decimal by 100, the decimal point is shifted to the left by two places. Then,

= 0.35



Thanks